# ODMixer: Fine-grained Spatial-temporal MLP for Metro Origin-Destination Prediction

### v2

@@ VI Conclusion
In this paper, we introduce ODMixer, a fine-grained spatial-temporal MLP architecture to address the metro OD prediction problem. Our ODMixer learns the short-term temporal relations of OD pairs by incorporating the Channel Mixer. The Multi-view Mixer efficiently captures OD pair relations from both origin and destination perspectives. With the integration of BTL, our ODMixer can perceive long-term traffic changes. Experimental results represent ODMixer’s outstanding performance on two large-scale datasets. Future directions for ODMixer involve incorporating additional city information, such as urban population distribution, regional composition, and Point of Interest (POI). Moreover, enhancing the model’s ability to learn the general pattern of traffic flow can improve its migration capability and scalability.

### v4

@@ VI Conclusion
In this paper, we introduce ODMixer, a fine-grained spatial-temporal MLP architecture for metro OD prediction problem. Our ODMixer learns the short-term temporal relations of OD pairs by incorporating the Channel Mixer. The Multi-view Mixer efficiently captures OD pair relations from both origin and destination perspectives. With the integration of BTL, our ODMixer can perceive long-term traffic changes. Experimental results represent ODMixer’s outstanding performance on two large-scale datasets. Future directions for ODMixer involve incorporating additional city information, such as urban population distribution, regional composition, and Point of Interest (POI). Moreover, enhancing the model’s ability to learn the general pattern of traffic flow can improve its migration capability and scalability. Deploying ODMixer in actual metro systems can facilitate the management and optimization of metro operations, thereby enhancing transportation efficiency.